OpenStudy (anonymous):

aolvw by the substitution method x = 4y -1 3x + 5y = 31

OpenStudy (anonymous):

3(4y-1)+5y=31 12y-3+5y=31 17y=34 y=17 x=4(17)-1 x=68-1 x=67

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I need help finishing another question will you help?